pooja195 (pooja195):

Margie is responsible for buying a week's supply of food and medication for the dogs and cats at a local shelter. The food and medication for each dog costs twice as much as those supplies for a cat. She needs to feed 164 cats and 24 dogs. Her budget is $4240. How much can Margie spend on each dog for food and medication? Can someone guide me through this i have a hard time with problems like these >_< @Directrix @Zarkon please help ;-;

OpenStudy (fibonaccichick666):

Are you able to set up the equation? What have you tried so far?

pooja195 (pooja195):

Lol idk im confused =/

OpenStudy (compassionate):

You need a System of Equations.

pooja195 (pooja195):

lol thats what i can't figure out this is confusing :(

Nnesha (nnesha):

How much can Margie spend on each dog for food u need to find cost for dog or food so dog = x food = y

OpenStudy (fibonaccichick666):

I pass this to nnesha

OpenStudy (fibonaccichick666):

She got it

pooja195 (pooja195):

i know about the dogs but not the food

pooja195 (pooja195):

NNesha thanks I got it! :*

Nnesha (nnesha):

wait lets start it again

Nnesha (nnesha):

nope !!!! okay let me know what you got show your work ??

pooja195 (pooja195):

164c + 24d = 4240 d = 2c 164c + 24(2c) = 4240 164c + 48c = 4240 212c = 4240 c = 20 She can spend $20 on each cat. d = 2c d = 2(20) = 40 She can spend $40 on each dog.

Nnesha (nnesha):

okay now start it again x = dog cats y = dogs cost